OpenStudy (anonymous):

each roller under a conveyor belt has a radius of 0.5 meters. the rollers turn at a rate of 30 revolutions per minute. what is the linear velocity of the conveyor belt

OpenStudy (anonymous):

this is physics =] \[30 revs*\frac{2pi}{1rev}*\frac{1min}{60s}=\frac{60\pi rads }{s}\] this is your angular velocity

OpenStudy (anonymous):

ooops it should be \[\frac{60\pi rads}{60s}=\pi rad/s\]

OpenStudy (anonymous):

so just 3.14 rad/s

OpenStudy (anonymous):

now what relationship is htere to go from angular velocity to

OpenStudy (anonymous):

linear velocity

OpenStudy (anonymous):

so what would the answer be and the units are m/s

OpenStudy (anonymous):

yes the relationship between linear and angular is \[r\omega=v\] you know the angular velocity and the radius

OpenStudy (anonymous):

Im confussed of what the final answer is

OpenStudy (mertsj):

If the radius is .5 then one revolution is pi since the circumference is pi times diameter. In 30 revolutions, it goes 30 pi or 3.14 times 30 or 94.2 meters per minute.

OpenStudy (anonymous):